OpenStudy (trisarahtops):

Please Help!!! I'll give a MEDAL to anyone who tries!! How do you create a fourth-degree polynomial with only two real zeros. Demonstrate how to do this and explain your steps.

jimthompson5910 (jim_thompson5910):

Those two zeros could repeat themselves twice. So for instance you could have the two zeros: 5,7 but since each repeats twice, you really have these 4 zeros: 5,5,7,7

jimthompson5910 (jim_thompson5910):

let me know if that's enough info to get started

OpenStudy (trisarahtops):

I don't think i understand? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you could easily have zeros repeat themselves

jimthompson5910 (jim_thompson5910):

so if you had the four zeros 5,5,7,7 this would mean that x = 5, x = 5, x = 7, x = 7 x-5=0, x-5=0, x-7=0, x-7=0 (x-5)(x-5)(x-7)(x-7) = 0 I'll let you expand that out

OpenStudy (trisarahtops):

so would it also work if like for example we had (x+2)(x+2)(x+2)(x-5)? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes that would work as well expand that out and tell me what you get